# Prediction-Enhanced Monte Carlo: A Machine Learning View on Control Variate

### v2

@@ 2. Prediction-Enhanced Monte Carlo
When $\boldsymbol{Y}$ can be simulated and $f$ evaluated, then standard MC amounts to generating many $f(\boldsymbol{Y}_{i})$, $i=1,\ldots,n$ and outputting their average. While this approach is demanding when $f$ and $\boldsymbol{Y}$ are expensive, it is unbiased and its statistical error can be quantified straightforwardly via standard confidence intervals.
@@
To reduce computation, one can approximate $\mu(\boldsymbol{\theta})$ with an ML regression model, say $\hat{\mu}(\boldsymbol{\theta})$, which can be evaluated more cheaply than simulating many $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})$’s. This natural approach, however, would face challenges at multiple levels when reliability and error quantification is important: From a theoretical standpoint, ML generalization bounds typically assess training errors only within the model class, i.e., $\hat{\mu}(\boldsymbol{\theta})-\mu^{*}(\boldsymbol{\theta})$ where $\mu^{*}(\cdot)$ is the best in-class model that can be different from $\mu(\cdot)$. Moreover, even assuming $\mu^{*}(\cdot)$ is a “rich" class, these bounds still only assess errors in an aggregate, worst-case fashion, i.e., high-probability bounds treating $\boldsymbol{\theta}$ as a random covariate. On the practical front, training-testing split or cross-validation succumbs to similar issues, in that it cannot measure bias against $\mu(\cdot)$, nor at the level of each $\boldsymbol{\theta}$. That is, in terms of reliability and uncertainty handling, MC appears superior by providing accurate, straightforward and instance-specific error assessment on each estimate of $\mu(\boldsymbol{\theta})$. None of these advantages are available for ML despite its computational edge.
@@
An idea to leverage the respective advantages of MC and ML is to use control variate (CV), a classical variance reduction method. CV aims to reduce the error of each MC run, consequently the required replication size, by utilizing information from “auxiliary” simulation outputs. Specifically, in generating $f(\boldsymbol{Y}_{i})$, suppose we can also generate auxiliary output $\boldsymbol{X}_{i}$ and evaluate $g(\boldsymbol{X}_{i})$ for some function $g$. Then we can use $f(\boldsymbol{Y}_{i})-g(\boldsymbol{X}_{i})+\mathbb{E}[g(\boldsymbol{X})]$ as an unbiased replicate for $\mathbb{E}[f(\boldsymbol{Y})]$. In particular, if $f(\boldsymbol{Y})$ and $g(\boldsymbol{X})$ are highly correlated, this replicate will have a smaller variance than $f(\boldsymbol{Y}_{i})$.
@@
Naturally, one can think of using ML to train a good predictor $g(\boldsymbol{X})$ for $f(\boldsymbol{Y})$, which would be highly correlated with $f(\boldsymbol{Y})$ and serve as an effective CV. However, taking aside the question of how to construct $\boldsymbol{X}$ and $g$, the CV method requires knowledge of $\mathbb{E}[g(\boldsymbol{X})]$, the closed-from mean on the auxiliary output. This knowledge is the key convertible into variance reduction. Yet, it is seldom available and has apparently severely limited the historical applicability of CV [61] and arguably even the more recent ML-inspired approaches [129, 121, 88].
We introduce Prediction-Enhanced Monte Carlo (PEMC), a framework that fulfills these objectives. PEMC leverages the classical control variate idea but generalizes its variance-reduction notion to a scheme-wide, modernized view, allowing for effective integration of MC with modern ML. Concretely, PEMC aims to estimate (1) by
@@
Here, $\boldsymbol{X}$ is an auxiliary variable, coupled with $\boldsymbol{Y}$, that serves as features for predicting $f(\boldsymbol{Y})$ and is also chosen to have a marginal distribution that is both efficiently simulatable and highly parallelizable, allowing large number $N$ independent samples of $\tilde{\boldsymbol{X}}$ with the same marginal as $\boldsymbol{X}$ to be generated at low cost. The function $g$ is a prediction model trained via squared-error loss using MC sample under $(\boldsymbol{X},\boldsymbol{Y})\sim\mathbb{P}_{\boldsymbol{\theta}}$ for various $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$, which aims to approximate the function $g\bigl{(}\boldsymbol{\theta},\boldsymbol{X}\bigr{)}\approx\mathbb{E}_{\boldsymbol{\theta}}\bigl{[}f_{\boldsymbol{\theta}}\bigl{(}\boldsymbol{Y}\bigr{)}\big{|}\boldsymbol{\theta},\boldsymbol{X}\bigr{]}$. As we will see, the selection of an appropriate $\boldsymbol{X}$ is a pivotal component of PEMC, analogous to feature selection in common ML tasks. The key lies in selecting $\boldsymbol{X}$ to be both a feature for predicting $f(\boldsymbol{Y})$ and an efficiently simulatable, parallelizable random variable. Finding a good $\boldsymbol{X}$ often requires domain knowledge. In particular, given the whole path $\boldsymbol{Y}$, one can choose $\boldsymbol{X}:=\phi(\boldsymbol{Y})$ to be some low-dimensional transformation $\phi$ of $\boldsymbol{Y}$ (e.g., ingredients for constructing $\boldsymbol{Y}$). For instance, for SDEs, one can define $\boldsymbol{X}$ as the sum of driving Brownian increments that generate $\boldsymbol{Y}$. This choice could capture a substantial portion of the SDE’s variance, making $\boldsymbol{X}$ a good feature for predicting $f(\boldsymbol{Y})$, while its Gaussianity allows it to be generated cheaply and in parallel. We explore this in detail in subsequent sections.
Many implementation details remain unspecified in the basic description above. For instance, how exactly to train g, how to select the sample sizes $n$ and $N$, and how to adapt PEMC to a conditional setting where $\theta\mapsto\boldsymbol{X}(\theta)$, $\boldsymbol{Y}(\theta)$, and $\mathbb{E}_{\theta}$ all depend on some parameter $\theta$. The next sections will address these questions in depth, providing both theoretical justification and practical guidelines for deploying PEMC in diverse scenarios.
3. Prediction-Enhanced Monte Carlo: Framework via Example
In this section, we employ Asian option pricing as a running example on how to apply PEMC. Asian options are path-dependent derivatives whose payoffs depend on the average price of the underlying asset over a specified period. As mentioned in [23] and [44], closed-form solutions for the pricing of arithmetic average Asian options are generally not available.
@@ 3.1. Problem Setup
For Asian option, we can specify the framework as:
@@
    Simulation Parameters $\boldsymbol{\theta}_{\text{simulation}}$: These parameters also specify $\mathbb{E}_{\text{risk neutral measure}}$. The distinction between $\boldsymbol{\theta}_{\text{simulation}}$ and $\boldsymbol{\theta}_{\text{model}}$ is somewhat artificial, primarily stemming from the convention that not all simulation parameters are calibrated externally, but rather serve as hyperparameters of the Monte Carlo simulation itself (either customized or directly observed). For example, when one simulates the Heston model above using Euler discretization scheme, [^1] As mentioned in [44], Euler scheme was the primary method for simulation of Heston model until the exact simulation of [31] is developed. However, due to its significant computational demands, the exact method is less commonly used in practice and simpler, biased simulation method such as Quadratic Exponential [3] or Euler scheme remains prevalent (see, e.g., [91]). we use $\boldsymbol{\theta}_{\text{simulation}}$ to specify the initial stock price $S_{0}$, the initial volatility $\nu_{0}$, the time horizon $T$, and discretization time step $\Delta t$. The choice of these parameters can impact both the accuracy and efficiency of the simulation. While not typically part of the model calibration process, these parameters also play a crucial role in describing risk-neutral measure.
@@
    Payoff Function Parameters $\boldsymbol{\theta_{\text{payoff}}}$: This parameter $\boldsymbol{\theta_{\text{payoff}}}$ parametrizes the payoff function $f_{\boldsymbol{\theta}_{\text{payoff}}}$. In options pricing, this specifies the details of the contract. For example, given the full price path $\boldsymbol{Y}:=\{S_{t}\}_{t}$, the payoff function for an arithmetic Asian options takes the form
    $P_{A}(\{S_{t}\}_{t})=\max\left(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K,0\right)$
    where $\boldsymbol{\theta_{\text{payoff}}}$ specifies the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$. Notice we have simply omitted the discount factor $e^{-rT}$. Note that the discrete sampling defied by the payoff does not necessarily determine the Euler discretization in time and in fact the latter is usually much more frequent for accuracy.
Remark Having specified the distinct roles of $\boldsymbol{\theta}_{\text{model}}$, $\boldsymbol{\theta}_{\text{simulation}}$ and $\boldsymbol{\theta}_{\text{payoff}}$, we will henceforth refer to them collectively as $\boldsymbol{\theta}$ when no ambiguity arises.
3.2. Prediction-Enhanced Monte Carlo: Motivation
As we have seen, Asian options are path-dependent derivatives whose payoffs depend on the average prices of the underlying asset. Generally, closed-form solutions for these options are unavailable, prompting numerous studies to propose various numerical methods for their pricing. In particular, the arithmetic Asian option under the geometric Brownian motion (GBM) model is one of the rare cases where a closed-form geometric Asian option price exists and can serve as a control variate. Beyond simple models like GBM, closed-form CV solutions are unavailable [44]. To address these challenges, the Prediction-Enhanced Monte Carlo (PEMC) framework aims to reduce the number of required Monte Carlo samples by generating a machine learning-based control variate that is both highly parallelizable and efficiently evaluable. Crucially, our framework preserves the unbiased nature of the Monte Carlo estimator, a property that [60] highlights as “fundamental to the wide applicability of Monte Carlo methods in finance,” achieving a remarkable elimination of any potential bias introduced by any black-box prediction model.
In PEMC, we do not specify the training algorithm or model architecture for this prediction model, as our framework is compatible with any black-box model. However, we want to find some input feature $\boldsymbol{X}:=\phi(\boldsymbol{Y})$ that is some typically low-dimensional transformation $\phi$ of $\boldsymbol{Y}$ with:
@@ 3.2. Prediction-Enhanced Monte Carlo: Motivation
The selection of an appropriate $\boldsymbol{X}$ is a pivotal component of our framework, analogous to feature selection in machine learning tasks. This process necessitates domain expertise. For now, we proceed under the assumption that a suitable $\boldsymbol{X}$ has been identified.
3.3. Prediction-Enhanced Monte Carlo: Prediction Model

### v3

@@ 2.1. Existing Challenges of MC, ML, and CV
When $\boldsymbol{Y}$ can be simulated and $f_{\boldsymbol{\theta}}$ evaluated, then standard MC amounts to generating many $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})$, $i=1,\ldots,n$ and outputting their average. This approach is unbiased and its statistical error can be quantified straightforwardly via standard confidence intervals. However, it is demanding when evaluating $f_{\boldsymbol{\theta}}$ and simulating $\boldsymbol{Y}$ is expensive.
@@
To reduce computation, one can approximate $\mu(\boldsymbol{\theta})$ with a ML regression model, say $\hat{\mu}(\boldsymbol{\theta})$, which can be evaluated more cheaply than simulating many $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})$’s. This natural approach, however, would face challenges at multiple levels when reliability and error quantification is important: From a theoretical standpoint, ML generalization bounds typically assess training errors only within the model class, i.e., $\hat{\mu}(\boldsymbol{\theta})-\mu^{*}(\boldsymbol{\theta})$ where $\mu^{*}(\cdot)$ is the best in-class model that can be different from $\mu(\cdot)$. Moreover, even assuming $\mu^{*}(\cdot)$ is a “rich" class, these bounds still only assess errors in an aggregate, worst-case fashion, i.e., high-probability bounds treating $\boldsymbol{\theta}$ as a random covariate. On the practical front, training-testing split or cross-validation succumbs to similar issues, in that it cannot measure bias against $\mu(\cdot)$, nor at the level of each $\boldsymbol{\theta}$. That is, in terms of reliability and uncertainty handling, MC appears superior by providing accurate, straightforward and instance-specific error assessment on each estimate of $\mu(\boldsymbol{\theta})$. None of these advantages are available for ML despite its computational edge.
@@
An idea to leverage the respective advantages of MC and ML is to use CV which, as mentioned in Section 1, is a classical variance reduction method. CV aims to reduce the error of each MC run, consequently the required replication size, by utilizing information from “auxiliary” simulation outputs. Specifically, in generating $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})$, suppose we can also generate auxiliary output $\boldsymbol{X}_{i}$ and evaluate $g(\boldsymbol{\theta},\boldsymbol{X}_{i})$ for some function $g$. Then we can use $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i})+\mathbb{E}[g(\boldsymbol{\theta},\boldsymbol{X})]$ as an unbiased replicate for $\mathbb{E}_{\boldsymbol{\theta}}[f_{\boldsymbol{\theta}}(\boldsymbol{Y})]$. In particular, if $f_{\boldsymbol{\theta}}(\boldsymbol{Y})$ and $g(\boldsymbol{\theta},\boldsymbol{X})$ are highly correlated, this replicate will have a smaller variance than $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})$.
@@
Naturally, one can think of using ML to train a good predictor $g(\boldsymbol{\theta},\boldsymbol{X})$ for $f_{\boldsymbol{\theta}}(\boldsymbol{Y})$, which would be highly correlated with $f_{\boldsymbol{\theta}}(\boldsymbol{Y})$ and serve as an effective CV. However, taking aside the question of how to construct $\boldsymbol{X}$ and $g$, the CV method requires knowledge of $\mathbb{E}_{\boldsymbol{\theta}}[g(\boldsymbol{\theta},\boldsymbol{X})]$, the closed-from mean on the auxiliary output. This knowledge is a key requirement for CV to achieve variance reduction. Yet, it is seldom available and has apparently severely limited the historical applicability of CV [62] and arguably even the more recent ML-inspired approaches [132, 124, 89].
2.2. Our Remedy and Requirements
Our PEMC aims to leverage the CV idea, with a twist to circumvent the need on the closed-form mean knowledge in order to allow for effective integration of MC with modern ML. This requires crucially a scheme-wide cost-aware view on variance reduction. Concretely, PEMC aims to estimate (1) by
@@ 2.2. Our Remedy and Requirements
Here, $\boldsymbol{X}$ is an auxiliary variable that is dependent on $\boldsymbol{Y}$. Suppose, for now, that we are given the $g$ function and the $\boldsymbol{X}$ variable. In equation (2), each pair of $(\boldsymbol{X}_{i},\boldsymbol{Y}_{i})$ is coupled and simulated together. On the other hand, each $\tilde{\boldsymbol{X}}_{j}$ is simulated independently from all $(\boldsymbol{X}_{i},\boldsymbol{Y}_{i})$’s, and has the same marginal distribution as $\boldsymbol{X}$. In this case, it is straightforward to see that (2) is an unbiased estimator of (1). On the other hand, in terms of cost efficiency, the total simulation cost consists of
$[\text{cost of generating each\ }f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i})]\times n+[\text{cost of generating each\ }g(\boldsymbol{\theta},\tilde{\boldsymbol{X}}_{j})]\times N$ (3)
while the total variance is
$\frac{Var(f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i}))}{n}+\frac{Var(g(\boldsymbol{\theta},\tilde{\boldsymbol{X}}_{j}))}{N}$ (4)
Suppose that $g$ and $\boldsymbol{X}$ are designed such that the cost of generating each $g(\boldsymbol{\theta},\tilde{\boldsymbol{X}}_{j})$ is relatively negligible compared to each original MC target replicate $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})$. Then, the total cost (3) will become approximately $[\text{cost of evaluating each\ }f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i})]\times n$, while the total variance (4) will become approximately $Var(f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i}))/n$. In this case, the cost-variance structure is the same as the conventional CV method: If $g(\boldsymbol{\theta},\boldsymbol{X})$ is a good predictor for $f_{\boldsymbol{\theta}}(\boldsymbol{Y})$, then $Var(f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i}))$ is smaller than $Var(f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i}))$ and thus we attain a variance reduction to the overall estimator. Like in the conventional CV method, to achieve a total reduction from $Var(f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i}))/n$ to $Var(f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i}))/n$, we generate $n$ samples of $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i})$, i.e., variance reduction is achieved per replication.
@@
While the above depicts the idealized situation where we assume away the cost of generating $g(\boldsymbol{\theta},\tilde{\boldsymbol{X}}_{j})$, it provides guidance on what we need practically to elicit a cost-variance structure in (2), as revealed by (3) and (4), that is more favorable than naive MC. First, we need $g(\boldsymbol{\theta},\boldsymbol{X})$ to serve as a good predictor for $f_{\boldsymbol{\theta}}(\boldsymbol{Y})$, but also that $\boldsymbol{X}$ has a marginal distribution that is both efficiently simulatable and highly parallelizable, allowing a large number $N$ of independent samples of $\tilde{\boldsymbol{X}}_{j}$ with the same marginal as $\boldsymbol{X}$ to be generated at low cost. To this end, the selection of an appropriate $\boldsymbol{X}$ is a pivotal component of PEMC, analogous to feature selection in common ML tasks but with the additional need of being an efficiently simulatable, parallelizable random variable. Finding a good $\boldsymbol{X}$ often requires domain knowledge. In particular, given the whole path $\boldsymbol{Y}$, one can choose $\boldsymbol{X}:=\phi(\boldsymbol{Y})$ to be some low-dimensional transformation $\phi$ of $\boldsymbol{Y}$ (e.g., ingredients for constructing $\boldsymbol{Y}$). For instance, for stochastic diferential equations (SDEs), one can define $\boldsymbol{X}$ as the sum of driving Brownian increments that generate $\boldsymbol{Y}$. This choice could capture a substantial portion of the SDE’s variance, making $\boldsymbol{X}$ a good feature for predicting $f(\boldsymbol{Y})$, while its Gaussianity allows it to be generated cheaply and in parallel.
Second, we require a function $g$ to convert $\boldsymbol{\theta}$ and $\boldsymbol{X}$ into a good predictor for $f_{\boldsymbol{\theta}}(\boldsymbol{Y})$. In this work, $g$ is pre-trained using a squared-error loss, with MC sample drawn under $(\boldsymbol{X},\boldsymbol{Y})\sim\mathbb{P}_{\boldsymbol{\theta}}$ for various $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$, which aims to obtain a function $g\bigl{(}\boldsymbol{\theta},\boldsymbol{X}\bigr{)}\approx\mathbb{E}_{\boldsymbol{\theta}}\bigl{[}f_{\boldsymbol{\theta}}\bigl{(}\boldsymbol{Y}\bigr{)}\big{|}\boldsymbol{\theta},\boldsymbol{X}\bigr{]}$. Note that this pre-training can be resource-consuming and takes significant amount of efforts. However, once this pre-training is done, the function $g$ and our approach can be used to improve the efficiency in evaluating $\mu(\boldsymbol{\theta})$ for any $\boldsymbol{\theta}$. For example, in the context of option pricing, the pre-trained $g$ can be stored as part of a model library tailored to a specific class of exotic options, enabling PEMC to be applied directly whenever new parameter configurations arise.
3. An Illustrative Case Study: Pricing Asian Options with PEMC
In this section, we employ Asian option pricing as a running example to explain the detailed steps and their implications in PEMC. Asian options are path-dependent derivatives whose payoffs depend on the average price of the underlying asset over a specified period. As mentioned in [23] and [45], closed-form solutions for pricing arithmetic average Asian options are generally not available. However, when the asset price dynamic follows a geometric Brownian motion (GBM), there exists an excellent CV candidate, using the geometric average instead of the arithmetic average, whose mean is precisely known via the Black-Scholes formula. This latter case often serves as the “textbook" example to demonstrate the power of CV [61]. As an important message, our case study here illustrates how PEMC can achieve a comparable level of improvements as the “textbook" CV, for evaluating arithmetic average Asian option price when the dynamics does not follow GBM. In other words, PEMC is capable to extend the scope of classical CV to potentially a much wider range of problems.
@@ 3.1. Problem Setup
where the risk neural measure is a probability measure where all financial securities are valued as if investors are indifferent to risk, and thus all assets are assumed to earn the risk-free rate [40]. [^1] Under this measure, prices of financial securities are the discounted expected values of their future payoffs. For Asian option, we can specify the framework as:
@@
    Simulation Parameters $\boldsymbol{\theta}_{\text{simulation}}$: These parameters also specify $\mathbb{E}_{\text{risk neutral measure}}$. The distinction between $\boldsymbol{\theta}_{\text{simulation}}$ and $\boldsymbol{\theta}_{\text{model}}$ is somewhat artificial, primarily stemming from the convention that not all simulation parameters are calibrated externally, but rather serve as hyperparameters of the MC simulation itself (either customized or directly observed). For example, when one simulates the Heston model above using Euler discretization scheme, [^2] As mentioned in [45], Euler scheme was the primary method for simulation of Heston model until the exact simulation of [31] is developed. However, due to its significant computational demands, the exact method is less commonly used in practice and simpler, biased simulation method such as Quadratic Exponential [3] or Euler scheme remains prevalent (see, e.g., [92]). we use $\boldsymbol{\theta}_{\text{simulation}}$ to specify the initial stock price $S_{0}$, the initial volatility $\nu_{0}$, the time horizon $T$, and discretization time step $\Delta t$. The choice of these parameters can impact both the accuracy and efficiency of the simulation. While not typically part of the model calibration process, these parameters also play a crucial role in describing risk-neutral measure.
@@
    Payoff Function Parameters $\boldsymbol{\theta_{\text{payoff}}}$: This parameter $\boldsymbol{\theta_{\text{payoff}}}$ parametrizes the payoff function $f_{\boldsymbol{\theta}_{\text{payoff}}}$. In option pricing, this specifies the details of the contract. For example, given the full price path $\boldsymbol{Y}:=\{S_{t}\}_{t}$, the payoff function for an arithmetic Asian options takes the form
    $P_{A}(\{S_{t}\}_{t})=\max\left(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K,0\right)$ (7)
    where $\boldsymbol{\theta_{\text{payoff}}}$ specifies the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$. Notice we have simply omitted the discount factor $e^{-rT}$. Note that the discrete sampling defined by the payoff does not necessarily determine the Euler discretization in time and in fact the latter is usually much more frequent for accuracy.
Having specified the distinct roles of $\boldsymbol{\theta}_{\text{model}}$, $\boldsymbol{\theta}_{\text{simulation}}$ and $\boldsymbol{\theta}_{\text{payoff}}$, we will henceforth refer to them collectively as $\boldsymbol{\theta}$ when no ambiguity arises.
3.2. Feature Engineering to Design $\Large{X}$
We now follow the general ideas laid out in Section 2.2, starting with the creation of the variable $\boldsymbol{X}$ to couple with $\boldsymbol{Y}$. As described in Section 2.2, we want to find some input feature $\boldsymbol{X}:=\phi(\boldsymbol{Y})$ that is some low-dimensional transformation $\phi$ of $\boldsymbol{Y}$ with:
@@ 3.2. Feature Engineering to Design $\Large{X}$
To this end, we take $\boldsymbol{X}$ as the sum of two correlated Brownian increments driving (1), motivated by two key properties:
  • Predictability: The sum of Brownian increments explains away a portion of the variance of the SDE.
More precisely, when sampling $\boldsymbol{Y}$, we collect $\boldsymbol{X}$ to be a two-dimensional vector consisting of
simulated during each step of the Euler scheme with discretization scale $\Delta t$. The marginal of $\boldsymbol{X}$ is simply a two-dimensional Gaussian: $\frac{\boldsymbol{X}}{\sqrt{T}}\sim\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$.
Using Brownian increments, while natural, is not the only choice. As with most ML tasks, feature engineering that leverages domain expertise can yield better results. We illustrate with another example, the pricing of a floating strike lookback call option, under the Heston model (1). Here, the option payoff is given by:
$f(\{S_{t}\}_{t})=\max(S_{T}-\min_{0\leq t\leq T}S_{t},0)$
The rationale behind the construction of $\min_{0\leq t\leq T}(r-\eta/2)t+\sqrt{\eta}W_{t}^{S}$ term is grounded in the domain knowledge of the Heston model. Specifically, since $\eta$ represents the mean-reverting level of variance process $\nu_{t}$, we may approximate $\nu_{t}\approx\eta$ (i.e., assume fixed volatility) and simplify the asset price process to a GBM:
This approximation leads to the expression $f(\{S_{t}\}_{t})\approx S_{0}e^{(r-\nu/2)T+\sqrt{\nu}[\boldsymbol{X}]_{1}}-S_{0}e^{[\boldsymbol{X}]_{2}}$, which could serve as a potentially effective approximation of $f(\{S_{t}\}_{t})$. This corresponds directly to Property 1 in the construction of $\boldsymbol{X}$ in 3.3. For Property 2, the marginal distribution of $\boldsymbol{X}$ permits efficient and parallelizable simulation. In particular, applying Girsanov’s Theorem enables one to derive a closed-form expression for the joint density of $([\boldsymbol{X}]_{1},[\boldsymbol{X}]_{2})$ [79]. This result makes it straightforward to sample these components directly.
3.3. Building Prediction Model